OpenStudy (anonymous):

Solve the equation 6n + 11 = -7 + 3n for n.

OpenStudy (anonymous):

N= -6

OpenStudy (mathstudent55):

\(6n + 11 = -7 + 3n \) Bring all terms with n to the left side by subtracting 3n from both sides: \(3n + 11 = -7\) Bring all numbers to the right side by subtracting 11 from both sides: \(3x = -18\) Divide both sides by 3 to get x by itself: \( \dfrac{3x}{3} = \dfrac{-18}{3} \) \(x = -6\) Solution: \(x = -6 \)